# TGDataset: Collecting and Exploring the Largest Telegram Channels Dataset

### v2

@@ Abstract.
Telegram is a widely adopted instant messaging platform. It has become worldwide popular because of its emphasis on security and privacy and its social network features like the channels—virtual rooms in which only the admins can post and broadcast messages to all its subscribers. Channels are used to deliver live updates (e.g., weather alerts) and content to a large audience (e.g., COVID-19 announcements) but, unfortunately, also for disseminating radical ideologies and coordinating attacks (e.g., Capitol Hill riot). This paper introduces the TGDataset, the most extensive publicly available collection of Telegram channels, comprising 120,979 channels and over 400 million messages. After outlining the data collection process, we provide a comprehensive dataset overview. Utilizing language detection, we identify the prevalent languages within the dataset. Subsequently, we explore English channels, employing topic modeling to analyze the covered subjects. Finally, we discuss some use cases in which our dataset can be extremely useful to better understand the Telegram ecosystem and to study the diffusion of questionable news. Alongside the raw dataset, we release the scripts used in our analysis, as well as a list of channels associated with a novel conspiracy theory called Sabmyk.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Information retrievalccs: Information systems Social networks
@@ 1. Introduction
In today’s digital age, instant messaging apps have become pervasive, and Telegram stands out as one of the leading platforms. Its emphasis on user privacy is one key factor for its increasing popularity. Indeed, the change in WhatsApp’s privacy policy in early 2021 prompted a significant migration of users to Telegram (for, 2021) for its commitment to user privacy. Yet, as is common with platforms prioritizing user privacy, its privacy-focused approach led malicious users to join the platform and use it for illegal activities, bringing Telegram to the center of several controversial incidents in recent years. Terrorists in Indonesia leveraged Telegram to spread radical ideologies and issue directives to conduct violent attacks (bbc, 2017). Neo-Nazis promoted their ideologies through Telegram channels and groups (tel, 2019). Crypto investors utilize Telegram channels to coordinate large-scale market manipulations, such as pump and dump schemes (La Morgia et al., 2020). Furthermore, channels featuring revenge porn (rev, 2020) and child pornography (chl, 2019) are unfortunately not uncommon. This highlights the complex nature of Telegram and its largely unknown darker aspects that deserve investigation.
In this study, we introduce the TGDataset, a collection of public Telegram channels of over 120,000 channels and 400 million messages, that we publicly release (Anonymous, 2024b). To the best of our knowledge, the TGDataset is the largest publicly available collection of Telegram channels. We begin by providing an overview of the dataset, quantifying the main features of the TGDataset. Then, we perform some vertical analysis, detecting the languages that are used the most in the channels and the topics covered by the English channels. Among many legitimate topics, we find channels publicizing porn, carding, violence, hacking, and white supremacy content. Studying the creation date of the channels to understand the evolution of Telegram over time, we find consistent growth in the creation of channels on Telegram, with Russian channels taking the lead until the ban in 2018, followed by English channels surpassing them in 2021, likely due to the policy changes in WhatsApp and the resulting user migration.
While previous datasets on Telegram focused on specific research objectives that limited their scope to channels related to a particular language, country, or topic, our dataset provides a more comprehensive representation of Telegram. Indeed, it includes channels dealing with various subjects in different languages. This aspect makes the TGDataset an excellent starting point for studying several phenomena occurring on Telegram. One of these phenomena concerns the political leaning of the platform and the sharing of questionable content, a subject widely explored on other Online Social Networks (OSNs) but not on Telegram. Moreover, the dataset includes numerous channels disseminating conspiracy theories, including newer ones (i.e., Sabmyk). Furthermore, the massive presence of channels running borderline activities (i.e., carding), calling to violence, or promoting Nazi ideologies suggests that the TGDataset represents a valuable resource for researchers willing to investigate and shed light on this dark side of Telegram.
@@ 2. Background: Telegram
Started in 2013, Telegram is now a popular instant messaging platform, with more than 500 million active users in 2021 (Tel, 2021). Telegram allows users to post content such as text messages, images, audio, stickers, videos, and files weighing up to 2 GB. In addition to private chats (one-to-one communication), Telegram provides two other solutions to communicate: Groups and Channels.
@@ 3. Related works
Several studies have been conducted on the Telegram platform and related issues. Hashemi et al. (Hashemi and Chahooki, 2019) collected Telegram Iranian channels and groups to distinguish high-quality groups (e.g., business groups) from low-quality groups (e.g., dating groups). They found that high-quality groups had longer messages and more user engagement. Nobari et al. (Dargahi Nobari et al., 2017) performed a structural and topical analysis of Telegram messages, building a dataset of over 2,000 groups and channels and constructing a graph based on mentions. This study found that the PageRank algorithm was not effective in identifying high-quality Telegram channels. Jalilvand et al. (Jalilvand and Neshati, 2020) tackled the challenge of ranking channels related to a user request on Telegram. Ng et al. (Ng and Yuan, 2020) analyzed a Singapore-based COVID-19 Telegram group with over 10,000 participants, examining group opinions over time. Their work revealed a peak in engagement when the Ministry of Health raised the disease alert level but a subsequent decline in participation.
Similar to our work, Baumgartner et al. (Baumgartner et al., 2020) created the Pushshift Telegram dataset, containing both channels and groups. They collected over 27,800 channels and 317 million messages from 2.2 million unique users updated to November 2019. Differently from Baumgartner et al. (Baumgartner et al., 2020), the TGDataset boasts a more recent dataset, with the message history of channels aligned to July 2022, thus almost three year of more data. Moreover, while the Pushshift dataset contains both channels and groups, the TGDaset is made of only channels. Thus, narrowing the comparison on the Telegram channels, the TGDataset includes a significantly higher number of channels (approximately four times more) and a greater volume of messages posted in channels (more than double). Tab.1 reports the exact number of messages and channels within the two datasets.
@@
Pushshift 27,801 192,044,689 November 2019
@@
Finally, another important difference between the two datasets, is about the seed choice. Indeed, the Pushshift dataset used as seeds only channels of right-wing extremist politics or cryptocurrencies related channels, while our dataset start the collection from seed channels covering heterogeneous topics (see Sec. 4.1). This choice lead to build a dataset that should better represents the status of the Telegram channels ecosystem.
@@ 4. Data collection
Previous Telegram datasets have been created with specific research goals in mind, resulting in a limited scope of channels related to a particular language, country, or topic. For example, (Hashemi and Chahooki, 2019) only includes Iranian channels,  (Hoseini et al., 2021) only contains 151 channels related to QAnon, and the PushShift dataset focuses on channels discussing right-wing extremism and cryptocurrency. In contrast, our dataset aims to provide a global snapshot of the Telegram ecosystem. To achieve this goal, we require a dataset that provides an up-to-date representation of Telegram and encompasses a broad range of popular and interconnected channels. This is the motivation behind the creation of the TGDataset.
@@ 4.1. Dataset construction
For each seed channel, we proceed with downloading all its messages by leveraging the Telethon APIs (tel, 2021) (Step (2)), an open-source Python tool that provides access to the official Telegram APIs. Specifically, each call to download messages from a channel is performed five seconds after the previous one to avoid flooding the Telegram service. Once we retrieved the data about the channel, we proceeded with the data extraction (Step (3)): in particular, we retained all the information obtained by Telegram API, excluding media files such as images, videos, or PDFs. The reason behind this decision is to avoid storing copyrighted or illegal content. Anyway, even for media-based messages, we collect the name, author, and ID of the media content using the Telegram API, and we infer the file format based on the extension of the file name. Next, we store the extracted information in MongoDB (ATLAS, 2024), a NoSQL database (Step (4)). Finally, we parse the channel messages to identify forwarded messages and their original authors (Step (5)). If the author represents a channel that we have never seen, then the channel is added to the list of new channels. Conversely, if the author is a user or represents a group, we ignore the author of the forwarded message. To iteratively expand the TGDataset, we repeat the above procedure of channel discovery (Steps 2, 3, 4, and 5) using the newly found channels as new seeds. We stopped the iteration process on July 31, 2022. Finally, to align our dataset, for each discovered channel, we download all the missing messages until July 31, 2022.
@@ 4.2. Accessing the TGDataset and FAIR principles
Findable. We publicly released the TGDataset (Anonymous, 2024a) on Zenodo (European Organization For Nuclear Research and OpenAIRE, 2013), an open repository managed by CERN.
Accessible The dataset is made of 121 files, each of which contains a maximum of 1,000 channels. To ease the retrieval of single channels from the dataset, we store them in alphabetical order. In addition, each file’s name describes the index of the channels it contains. Given the large dimension of the dataset, even if compressed (approx. 71GB), we divided the dataset into four parts. In this way, it is possible to download and explore also a single portion of the dataset. Moreover, we released on a public GitHub repository (Anonymous, 2024b) a sample of the dataset. This enables users to explore the dataset’s structure without the need to download any portion of it.
@@
Reusable. We have outlined the methodology employed for data collection, and the data is open for anyone to use. We also released on the public GitHub repository (Anonymous, 2024b), Python scripts to load the data into a MongoDB (Banker et al., 2016) database, as well as scripts that show how to perform basic queries on the database (such as inserting new channels or retrieving channels by username or ID). Furthermore, these scripts can also replicate some of the analysis that we performed in this work, including Language Detection of the channels (Section 5.1) and Topic Modeling (Section 5.2). Finally, we released as CSV files the list of channels belonging to the Sabmyk network and the list of other channels promoting conspiracy theories (Sec. 6), the list of channels annotated with the detected writing language, and the one containing the mapping of inferred topics.
@@ 4.3. Dataset overview
To provide a general characterization of the channels contained in the TGDataset, in Fig. 3, we show the CDFs of the number of subscribers, posts published, and the ratio of forwarded messages for verified, scam, and standard channels. As shown in Fig. 3(a), verified channels (green line), in general, have more subscribers than scam (dashed orange line) and standard channels (dotted blue line). Indeed, the verified channels with more than 1,000 subscribers are 98.32%, while the scams and standard are 68.85% and 38.11%, respectively. Concerning the number of messages published, Fig. 3(b) suggests that verified channels tend to post more content than scams and standard ones. The standard channels with at least 1,000 messages are 57.16%, those verified are 83.79%, and the scams are 50.82%. Finally, we report the extent of use of the forwarding function among channels of the TGDataset. Fig. 3(c) reveals message forwarding is more prevalent in standard and scam channels. The standard channels with at least a 0.1 ratio of their messages that are forwarded are 33.5%, whereas the scams and the verified ones are 32.24% and 13.15%, respectively.
@@ 5.2. Topic modeling
The best UMass value (-0.69) is obtained with 13 topics. Tab. 3 shows the inferred topics and the top 10 keywords for each of them. Next, we group the channels according to the related topics using LDA. Tab. 5 lists the topics discovered in the TGDataset and the number of channels for each. These emerged topics are quite different from those covered by the seed channels (see Sec. 4). Indeed, some disappeared (e.g., Sport, Marketing), while other interesting topics came up. One of these is related to carding, the practice of selling full details of stolen credit cards or selling prepaid cards or other goods purchased with them. Similarly to what happens in dark web forums (Kigerl, 2020), carders (the people who own the stolen credit cards) use Telegram channels to place gift cards or goods for sale. In the TGDataset, we find 1,489 channels (7.53% of the English part of the dataset) offering this service. In particular, Telegram marked 45 carding channels as a scam suggesting that some of them do not deliver the offered service.
@@ 6. Use cases
The TGDataset has already been utilized in various studies (La Morgia et al., 2023; Imperati et al., 2023). For instance, La Morgia et al. (La Morgia et al., 2023) analyzed the problem of fake accounts on Telegram using the TGDataset, while Imperati et al. (Imperati et al., 2023) employed the TGDataset to identify 17,829 channels disseminating conspiracy-related content. In the following, this section describes other scenarios that can be explored and analyzed using the TGDataset.
@@
This analysis reveals a significant presence of Telegram channels that disseminate news from unreliable sources. Thus, the TGDataset could be a valuable resource for researchers aiming to understand the spread of misinformation, fake news, and conspiracy theories.
@@
Then, we use the Leiden algorithm (Traag et al., 2019), an algorithm for community detection that improves the Louvain algorithm (Ghosh et al., 2018). This approach finds a partition with 311 communities, 47 of which with more than one node. One of the communities, which consists of 236 channels and where English is the common language, has the peculiarity that all posts and messages forwarded from one channel to the other are about a new conspiracy theory called Sabmyk. Sabmik is a conspiracy theory that proposes itself as a better alternative to QAnon. It promotes a quasi-religion centered around a messianic figure known as Sabmyk (Sommerlad, 2021). This community of channels has, collectively, more than 1 million subscribers. The most popular channel in the community is Great Awakening Channel with 119,103 subscribers. According to our data, most of these channels were created at the beginning of 2021—130 of them (55.08%) between January and February—with the latest in our dataset created in February 2022. The nature of these channels is very different: Some of them are fake channels of celebrities (e.g., Mel Gibson, Keanu Reeves, Kanye West), while others target news outlets or official channels of national bodies (e.g., U. S. Marines Channel, U. S. Navy Channel) and others conspiracy theories like QAnon. Regardless of the name of the channel in the community, the content is always the same. Indeed, they recycle and forward messages among them: Out of the 1,203,986 messages published by these 236 channels, only 65,602 are unique. With this analysis we scratch the surface of the Sabmyk network. Still, more investigation are needed to understand how Sabmyk lures such a large number of subscribers, its ultimate goal, and its connection to other conspiracy theories.
@@
Carding and underground markets. In Section 5.2, we examine the main topics within our dataset. Upon investigating the channels associated with these topics, we discovered several engaging in borderline activities, such as running underground marketplaces. These channels sell electronics, Netflix accounts, and hacking tools at low prices. These types of marketplaces are often associated with carding. Consequently, it is worth delving further into these Telegram channels to determine the types of goods being sold, their origins, whether the channels are just an attempt to fraud subscribers or if they are actually engaged in more illegal activities, and whether there is any correlation with Dark Web marketplaces.
@@
Copyright Infringement and personal content. Recently, Telegram hit the news several times for its use for the distribution of copyrighted content (e.g., movies and software). For instance, Italian authorities seized 545 channels for copyright infringement (Maxwell, 2022). However, the issue extends beyond copyright violations. Indeed, the platform is frequently misused to distribute personal content without consent, such as the unauthorized distribution and sale of an Indian teacher’s course material (Singh, 2022). Even more troubling, Telegram is also used to share leaked nude photos (Lucy Swinnen and Gragnani, 2022) or facilitate revenge porn (Kraus, 2020). Joining the information about the topic of the channels, and the description of the messages removed by the platform (Tab. 4), we can observe that the four categories in which is most present the diffusion of copyrighted material are: Carding (448 channels), Indian Edu (219 channels), Video-game modding (208 channels), Adult content (157 channels). Starting from these channels, it is possible to study the diffusion of the phenomenon, understand how these channels are organized, and how they monetize.
@@
Cryptocurrencies frauds. We also found several (563) channels related to the cryptocurrency world. It is well known that Telegram channels are exploited by Pump and Dump groups (Xu and Livshits, 2019; Morgia et al., 2021) that perpetrate frauds on the crypto-market. We notice that some of the Pump and Dump channels monitored in (La Morgia et al., 2020) are also in our dataset. Previous works monitored the channels in a vertical disjoint way, focusing on the fraud and the groups’ mechanics. Thus, the TGDatast can help understand how these channels are connected, if they operate together, how they promote their services, and likely discover new channels that carry out similar activities.

### v3

@@ Abstract.
Telegram is a widely adopted instant messaging platform. It has become worldwide popular because of its emphasis on privacy and its social network features such as channels—virtual rooms in which only the admins can post and broadcast messages to all the subscribers. Channels are used to deliver live updates (e.g., weather alerts) and content to a large audience (e.g., COVID-19 announcements) but unfortunately also to disseminate radical ideologies and coordinate attacks such as the Capitol Hill riot.
This paper introduces the TGDataset, the most extensive publicly available collection of Telegram channels, comprising 120,979 channels and over 400 million messages. We outline the data collection process and provide a comprehensive overview of the data set. Using language detection, we identify the predominant languages within the dataset. We then focus on English channels, employing topic modeling to analyze the subjects they cover. Finally, we discuss some use cases in which our dataset can be instrumental in understanding the Telegram ecosystem and studying the diffusion of questionable news. Alongside the raw dataset, we release the scripts used in our analysis, as well as a list of channels associated with a novel conspiracy theory known as Sabmyk.
@@
journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709397isbn: 979-8-4007-1245-6/25/08ccs: Information systems Information retrievalccs: Information systems Social networks
[…]
@@ 1. Introduction
In today’s digital age, instant messaging apps have become ubiquitous, and Telegram stands out as one of the leading platforms. Its focus on user privacy is a significant factor behind its growing popularity (for, 2021). However, as is often the case with platforms that prioritize user privacy, this approach has also attracted malicious users who exploit the platform for illegal activities. While Telegram serves legitimate purposes, such as disseminating weather alerts or government updates (Leong, 2020), it is also misused for spreading radical ideologies (tel, 2019), orchestrating violent attacks (bbc, 2017), and engaging in market manipulations, such as pump-and-dump schemes (La Morgia et al., 2020). These dual uses underscore Telegram’s complex nature and highlight the need for a deeper exploration of its ecosystem.
In this study, we introduce and publicly release (La Morgia et al., 2024b) the TGDataset, a collection of public Telegram channels of over 120,000 channels and 400 million messages. To the best of our knowledge, the TGDataset surpasses existing datasets in scale and scope, offering an unprecedented opportunity to study the Telegram ecosystem comprehensively. Unlike prior datasets focused on specific topics or geographic regions, the TGDataset includes diverse channels covering multiple languages and subjects, enabling a more holistic understanding of the platform. In particular, the TGDataset enables the study of the platform’s political leanings and the sharing of questionable content, topics widely explored on other Online Social Networks (OSNs) but not on Telegram. It also includes numerous channels spreading conspiracy theories, as well as those engaged in borderline or illegal activities like carding, inciting violence, and promoting Nazi ideologies. Thus, with its scale and diversity, the TGDataset represents a valuable resource for researchers willing to investigate and shed light on both the legitimate and problematic aspects of Telegram’s ecosystem.
The main contributions of this work are the following:
  • We present the largest publicly available dataset of Telegram channels, enabling researchers to explore various phenomena on the platform. We perform analyses such as language detection and topic modeling to characterize the dataset’s contents, revealing dominant languages, prevalent topics, and temporal trends in channel creation.
  • We explore several potential use cases of the TGDataset. One application involves analyzing the diffusion of questionable content and biases across Telegram channels, providing insights into content moderation and information reliability. Another use case focuses on investigating networks that promote conspiracy theories, including emerging phenomena such as the Sabmyk network. Additionally, the dataset can be used to investigate channels engaged in borderline or illegal activities, such as carding, copyright infringement, and extremist propaganda.
  • Alongside the dataset, we release scripts for analysis, language detection, and topic modeling, fostering reproducibility and facilitating further research.
[…]
A figure showing three screenshots of a Telegram channel interface: (A) displays the channel’s header with four labeled elements - the title, subscriber count, description, and username; (B) shows the feed of posts published by the channel; (C) shows an example of a forwarded message within the channel. Refer to caption
@@ 2. Background: Telegram
Started in 2013, Telegram is now a popular instant messaging platform, with more than 500 million active users in 2021 (Tel, 2021). Telegram allows users to post content such as text messages, images, audio, stickers, videos, and files weighing up to 2 GB. In addition to private chats (one-to-one communication), Telegram provides two other ways to communicate: Through groups and channels.
@@ 3. Related works
Several studies have been conducted on the Telegram platform and related issues. Hashemi et al. (Hashemi and Chahooki, 2019) collected Telegram Iranian channels and groups to distinguish high-quality groups (e.g., business groups) from low-quality groups (e.g., dating groups). They found that high-quality groups had longer messages and more user engagement. Jalilvand et al. (Jalilvand and Neshati, 2020) tackled the challenge of ranking channels related to a user request on Telegram. Ng et al. (Ng and Yuan, 2020) analyzed a Singapore-based COVID-19 Telegram group with over 10,000 participants, examining group opinions over time. Their work revealed a peak in engagement when the Ministry of Health raised the disease alert level but a subsequent decline in participation. Nobari et al. (Dargahi Nobari et al., 2017) performed a structural and topical analysis of Telegram messages, building a dataset of over 2,500 channels and 54 groups and constructing a graph based on mentions. This study found that the PageRank algorithm was not effective in identifying high-quality Telegram channels. Analyzing the dataset released by Nobari et al., the TGDataset contains more than ten times the number of channels and is significantly more recent (July 2022 versus October 2016).
Similar to our work, Baumgartner et al. (Baumgartner et al., 2020) created the Pushshift Telegram dataset, containing both channels and groups. They collected over 27,800 channels and 317 million messages from 2.2 million unique users updated to November 2019. Differently from (Baumgartner et al., 2020), the TGDataset boasts a more recent dataset, with the message history of channels aligned to July 2022, thus almost three years of more data. Moreover, while the Pushshift dataset contains both channels and groups, the TGDaset is made of only channels. Thus, narrowing the comparison on the Telegram channels, the TGDataset includes a significantly higher number of channels (approximately four times more) and a greater volume of messages posted in channels (more than double). Table 1 provides the exact number of channels and messages contained in the dataset released by (Dargahi Nobari et al., 2017), Pushshift, and TGDataset.
@@
Nobari et al. (Dargahi Nobari et al., 2017) 2,556 36,928 October 2016
Pushshift (Baumgartner et al., 2020) 27,801 192,044,689 November 2019
@@
Finally, another important difference between the two datasets is about the seed choice. Indeed, the Pushshift dataset used as seed channels only those of right-wing extremist politics or cryptocurrencies-related channels, while our dataset starts the collection from seed channels covering heterogeneous topics (see Sec. 4.1). This choice led to building a dataset that should better represent the status of the Telegram channels ecosystem.
@@ 4. Data collection
Previous Telegram datasets have been created with specific research goals in mind, resulting in a limited scope of channels related to a particular language, country, or topic. For example, (Hashemi and Chahooki, 2019) only includes Iranian channels,  (Hoseini et al., 2023) only contains 151 channels related to QAnon, and the PushShift dataset focuses on channels discussing right-wing extremism and cryptocurrency. In contrast, our dataset aims to provide a global snapshot of the Telegram ecosystem. To achieve this goal, we require a dataset that provides an up-to-date representation of Telegram and encompasses a broad range of popular and interconnected channels. This is the motivation behind the creation of the TGDataset.
@@ 4.1. Dataset construction
[…]
A flow diagram illustrating the data collection process. The diagram shows the sequential steps involved in collecting, processing, and storing data from Telegram channels. Refer to caption
@@
For each seed channel, we proceed with downloading all its messages by leveraging the Telethon APIs (tel, 2021) (Step (2)), an open-source Python tool that provides access to the official Telegram APIs. Specifically, each call to download messages from a channel is performed five seconds after the previous one to avoid flooding the Telegram service. Once we retrieved the data about the channel, we proceeded with the data extraction (Step (3)): in particular, we retained all the information obtained by Telegram API, excluding media files such as images, videos, or PDFs. The reason behind this decision is to avoid storing copyrighted or illegal content. Anyway, even for media-based messages, we collect the name, author, and ID of the media content using the Telegram API, and we infer the file format based on the extension of the file name. Next, we store the extracted information in MongoDB (MONGODB, 2024), a NoSQL database (Step (4)). Finally, we parse the channel messages to identify forwarded messages and their original authors (Step (5)). If the author represents a channel that we have never seen, then the channel is added to the list of new channels. Conversely, if the author is a user or represents a group, we ignore the author of the forwarded message. To iteratively expand the TGDataset, we repeat the above procedure of channel discovery (Steps 2, 3, 4, and 5) using the newly found channels as new seeds. We stopped the iteration process on July 31, 2022. Finally, to align our dataset, for each discovered channel, we download all the missing messages until July 31, 2022.
@@ 4.2. Accessing the TGDataset and FAIR principles
Findable. We publicly released the TGDataset (La Morgia et al., 2024a) on Zenodo (European Organization For Nuclear Research and OpenAIRE, 2013), an open repository managed by CERN.
Accessible. The dataset is made of 121 files, each of which contains a maximum of 1,000 channels. To ease the retrieval of single channels from the dataset, we store them in alphabetical order. In addition, each file’s name describes the index of the channels it contains. Given the large dimension of the dataset, even if compressed (approx. 71GB), we divided the dataset into four parts. In this way, it is possible to download and explore also a single portion of the dataset. Moreover, we released a sample of the dataset on a public GitHub repository (La Morgia et al., 2024b). This enables users to explore the dataset’s structure without the need to download any portion of it.
@@
Reusable. We have outlined the methodology employed for data collection, and the data is open for anyone to use. We also released on the public GitHub repository (La Morgia et al., 2024b), Python scripts to load the data into a MongoDB (Banker et al., 2016) database, as well as scripts that show how to perform basic queries on the database (such as inserting new channels or retrieving channels by username or ID). Furthermore, these scripts can also replicate some of the analysis that we performed in this work, including Language Detection of the channels (Section 5.1) and Topic Modeling (Section 5.2). Finally, we released as CSV files the list of channels belonging to the Sabmyk network and the list of other channels promoting conspiracy theories (Sec. 6), the list of channels annotated with the detected writing language, and the one containing the mapping of inferred topics.
@@ 4.3. Dataset overview
[…]
A figure containing three CDF plots comparing scam, verified, and standard Telegram channels. The first plot shows the distribution of subscriber counts, the second shows the distribution of total messages posted, and the third shows the distribution of forwarded message ratios. These plots help visualize the differences in behavior and characteristics between the different types of channels.
@@
To provide a general characterization of the channels contained in the TGDataset, in Fig. 3, we show the CDFs of the number of subscribers, posts published, and the ratio of forwarded messages for verified, scam, and standard channels. As shown in Fig. 3(a), verified channels (green line), in general, have more subscribers than scam (dashed orange line) and standard channels (dotted blue line). Indeed, the verified channels with more than 1,000 subscribers are 98.32%, while the scams and standard are 68.85% and 38.11%, respectively. Concerning the number of messages published, Fig. 3(b) suggests that verified channels tend to post more content than scams and standard ones. The standard channels with at least 1,000 messages are 57.16%, those verified are 83.79%, and the scams are 50.82%. Moreover, we report the extent of use of the forwarding function among channels of the TGDataset. Fig. 3(c) reveals message forwarding is more prevalent in standard and scam channels. The standard channels with at least a 0.1 ratio of their messages that are forwarded are 33.5%, whereas the scams and the verified ones are 32.24% and 13.15%, respectively.
@@ 5.2. Topic modeling
The best UMass value (-0.69) is obtained with 13 topics. Tab. 3 shows the inferred topics and the top 10 keywords for each of them. Next, we group the channels according to the related topics using LDA. Tab. 5 lists the topics discovered in the TGDataset and the number of channels for each. These emerged topics are quite different from those covered by the seed channels (see Sec. 4). Indeed, some disappeared (e.g., Psychology, Marketing), while other interesting topics came up. This shift in topic distribution is due to the snowball approach, which adds new channels linked by message forwarding to the dataset. As a result, channels from larger, more interconnected networks with higher forwarding activity are more likely to be included. This method naturally shifts the dataset toward more viral or widely shared topics.
Among the new topics, one is related to carding, the practice of selling full details of stolen credit cards or selling prepaid cards or other goods purchased with them. Similarly to what happens in dark web forums (Kigerl, 2020), carders (the people who own the stolen credit cards) use Telegram channels to place gift cards or goods for sale. In the TGDataset, we find 1,489 channels (7.53% of the English part of the dataset) offering this service. In particular, Telegram marked 45 carding channels as scams, suggesting that some of them do not deliver the service they offer.
@@ 5.3. Temporal aspect
[…]
A time series plot showing the creation dates of Telegram channels over time, broken down into Russian, English, and all channels. The plot includes two significant events marked with dots: a green dot indicating when Russia banned Telegram, and an azure dot marking WhatsApp’s privacy policy announcement. Refer to caption
@@ 6. Use cases
The TGDataset has already been utilized in various studies (La Morgia et al., 2023, 2024c; Imperati et al., 2025). For instance, La Morgia et al. (La Morgia et al., 2023) analyzed the problem of fake accounts on Telegram using the TGDataset, while Imperati et al. (Imperati et al., 2025) employed the TGDataset to identify 17,829 channels disseminating conspiracy-related content. In the following, this section describes other scenarios that can be explored and analyzed using the TGDataset.
@@
This analysis reveals a significant presence of Telegram channels disseminating news from unreliable sources. Thus, the TGDataset could be a valuable resource for researchers aiming to understand the spread of misinformation, fake news, and conspiracy theories. This use case involves using the content of the text messages posted by the channels (represented by the Message field) to extract links for further processing, and the forwarding information, including whether the message was forwarded and its source (contained in the Is Forwarded and Forwarded From fields).
@@
Then, we use the Leiden algorithm (Traag et al., 2019), an algorithm for community detection that improves the Louvain algorithm (Ghosh et al., 2018). To ensure the accuracy of our partition, we used modularity, a validation metric that measures how much better our partitioning is compared to a random partition. Thus, we compute the optimal number of communities with respect to modularity, achieving a high score of 0.78, which indicates a strong community structure remarkably better than random partitioning. This approach finds a partition with 311 communities, 47 of which with more than one node. One of the communities, which consists of 236 channels and where English is the common language, has the peculiarity that all posts and messages forwarded from one channel to the other are about a new conspiracy theory called Sabmyk. Sabmik is a conspiracy theory that proposes itself as a better alternative to QAnon. It promotes a quasi-religion centered around a messianic figure known as Sabmyk (Sommerlad, 2021). This community of channels has, collectively, more than 1 million subscribers. The most popular channel in the community is Great Awakening Channel with 119,103 subscribers. According to our data, most of these channels were created at the beginning of 2021—130 of them (55.08%) between January and February—with the latest in our dataset created in February 2022. The nature of these channels is very different: Some of them are fake channels of celebrities (e.g., Mel Gibson, Keanu Reeves, Kanye West), while others target news outlets or official channels of national bodies (e.g., U. S. Marines Channel, U. S. Navy Channel) and others conspiracy theories like QAnon. Regardless of the name of the channel in the community, the content is always the same. Indeed, they recycle and forward messages among them: Out of the 1,203,986 messages published by these 236 channels, only 65,602 are unique. With this analysis we scratch the surface of the Sabmyk network. Still, more investigation are needed to understand how Sabmyk lures such a large number of subscribers, its ultimate goal, and its connection to other conspiracy theories.
Performing this study requires several key elements of the dataset. By analyzing the content of the Message field, it is possible to identify the topics within conspiracy theories. With the Is Forwarded field, we can understand if the message has been forwarded, and in the case it is, which was the source channel leveraging the Forwarded From field. This information makes it possible to build a graph representing connections between channels and to detect the communities as we did for our analysis.
@@
Carding and underground markets. In Section 5.2, we examine the main topics within our dataset. Upon investigating the channels associated with these topics, we discovered several engaging in borderline activities, such as running underground marketplaces. These channels sell electronics, Netflix accounts, and hacking tools for low prices. These types of marketplaces are often associated with carding. Consequently, it is worth delving further into these Telegram channels to determine the types of goods being sold, their origins, whether the channels are just an attempt to fraud subscribers or if they are actually engaged in more illegal activities, and whether there is any correlation with Dark Web marketplaces.
For this analysis, the Message field is crucial for identifying the types of goods sold within the channels, such as electronics, Netflix accounts, and hacking tools. Analyzing the content of the messages and extracting the URLs makes it possible to determine whether the channels are involved in fraudulent activities or illegal transactions.
@@
Copyright Infringement and personal content. Recently, Telegram hit the news several times for its use for the distribution of copyrighted content (e.g., movies and software). For instance, Italian authorities seized 545 channels for copyright infringement (Maxwell, 2022). However, the issue extends beyond copyright violations. Indeed, the platform is frequently misused to distribute personal content without consent, such as the unauthorized distribution and sale of an Indian teacher’s course material (Singh, 2022). Even more troubling, Telegram is also used to share leaked nude photos (Lucy Swinnen and Gragnani, 2022) or facilitate revenge porn (Kraus, 2020). Joining the information about the topic of the channels, and the description of the messages removed by the platform (Tab. 4), we can observe that the four categories in which is most present the diffusion of copyrighted material are: Carding (448 channels), Indian Edu (219 channels), Video-game modding (208 channels), Adult content (157 channels). Starting from these channels, it is possible to study the diffusion of the phenomenon, understand how these channels are organized, and how they monetize. For this analysis, the Message field is essential for identifying illicit content, such as copyrighted material, while the Description field could provide additional context about channel activities.
@@
Cryptocurrencies frauds. We also found several (563) channels related to the cryptocurrency world. It is well known that Telegram channels are exploited by Pump and Dump groups (Xu and Livshits, 2019; Morgia et al., 2021) that perpetrate frauds on the crypto-market. We notice that some of the Pump and Dump channels monitored in (La Morgia et al., 2020) are also in our dataset. Previous works monitored the channels vertically, focusing on the fraud and the groups’ mechanics. Thus, using the Message, Forwarded from, and Description fields, the TGDatast can help understand how these channels are connected, if they operate together, how they promote their services, and likely discover new channels that carry out similar activities.